Parker Plastic, Inc., manufactures plastic mats to use with rolling office chairs. Its standard cost information for last year follows:

Standard Quantity Standard Price Standard Unit Cost
Direct materials (Plastic) 10 Sq Ft 0.82 per square feet 8.20
Direct Labor 0.25 hr 10.20 per hours 2.55
VMOH 0.25 hr 1.40 per hour 0.35

Fixed MOH 0.30

Number of Units Produced and Sold 1,020,000
Number of square feet of Plastic used 11,000,000
Cost of Plastic Purchased and used 8,800,000
Number of Labor Hours Worked 300,000
Direct Labor Cost 3,000,000
Variable overhead Cost 400,000
Fixed Overhead Cost 357,000

1. Calculate Parker Plastic’s variable overhead rate and efficiency variances and its over- or underapplied variable overhead. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)
Answer #1
Variable overhead rate variance 20000 F =400000-(300000*1.4)
Variable overhead efficiency variance 63000 U =1.4*(300000-1020000*0.25)
Variable overhead spending variance 43000 U =400000-(1020000*0.35)